# Course Allocation with Credits via Stable Matching

### v1

@@ 5.1 No lower quotas
For every $w_{i}\in W$, we create the following students: $s_{i}^{1}$, $s_{i}^{2}$, and $s_{i}^{3}$, with $T(s_{i}^{1})=T(s_{i}^{2})=T(s_{i}^{3})=2$. We also create courses $d_{i}^{1}$ and $d_{i}^{2}$, with $O(d_{i}^{1})=O(d_{i}^{2})=2$. Students $s_{i}^{1}$, $s_{i}^{2}$ and $s_{i}^{3}$ belong to subset $S_{i}\subseteq S$, and courses $d_{i}^{1}$ and $d_{i}^{2}$ belong to subset $D_{i}\subseteq C$. For every $w_{i}\in W$, $s_{i}^{1}\in S^{1}\subseteq S$, $s_{i}^{2}\in S^{2}\subseteq S$, and $s_{i}^{3}\in S^{3}\subseteq S$. Similarly, $d_{i}^{1}\in D^{1}\subseteq C$, and $d_{i}^{2}\in D^{2}\subseteq C$.
@@
For every $u_{j}\in U$, we create a course $c_{j}$, with $O(c_{j})=1$. Additionally, we create three courses $e_{1}$, $e_{2}$ and $f$, with $O(e_{1})=O(e_{2})=2$ and $O(f)=1$, as well as $n_{1}-K$ students labelled $p_{1}$ to $p_{n_{1}-K}$, such that, for every $p_{k}$, $T(p_{k})=1$. The upper quota of every course is $1$, except for $e_{1}$ and $e_{2}$, which have upper quota $n_{2}-K$, and $f$, which has upper quota $K$. For every $u_{j}\in U$, $c_{j}\in C^{\prime}\subseteq C$. Every $p_{k}$ belongs to $P\subseteq S$.
The master list of students (MLS) is
$s_{1}^{1}\succ_{MLS}s_{1}^{2}\succ_{MLS}s_{2}^{1}\succ_{MLS}s_{2}^{2}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{1}\succ_{MLS}s_{n_{2}}^{2}\succ_{MLS}\\ \succ_{MLS}s_{1}^{3}\succ_{MLS}s_{2}^{3}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{3}\succ_{MLS}p_{1}\succ_{MLS}p_{2}\succ_{MLS}...\succ_{MLS}p_{n_{1}-K}$ (2)
@@
That is, the first students belong to subsets $S^{1}$ and $S^{2}$, such that every $s_{i}^{1}$ is followed by $s_{i}^{2}$. Then we place students from $S^{3}$, and finally students from $P$, in all cases in increasing subscript order. The master list of courses (MLC) is
$c_{1}\succ_{MLC}c_{2}\succ_{MLC}...\succ_{MLC}c_{n_{1}}\succ_{MLC}d_{1}^{1}\succ_{MLC}d_{1}^{2}\succ_{MLC}d_{2}^{1}\succ_{MLC}d_{2}^{2}\succ_{MLC}...\succ_{MLC}\\ \succ_{MLC}d_{n_{2}}^{1}\succ_{MLC}d_{n^{2}}^{2}\succ_{MLC}e_{1}\succ_{MLC}e_{2}\succ_{MLC}f$ (3)
@@
If in the Exact-MM instance $w_{i}$ is matched with $u_{j}$, then $s_{i}^{1}$ is to be matched with $c_{j}$, $s_{i}^{2}$ with $d_{i}^{1}$ and $s_{i}^{3}$ with $d_{i}^{2}$. If, however, $w_{i}$ is unmatched, $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{2}$, and $s_{i}^{3}$ with $e_{1}$ and $e_{2}$. Student $s_{i}^{3}$ contains courses $c_{j}$ and $c_{k}$ in her preference list to make sure that a course-complete first-coalition-stable matching in $I$ corresponds to a maximal matching in $G$, so that if $s_{i}^{3}$ is matched with $e_{1}$ and $e_{2}$, some students from $S^{1}$ or $S^{2}$ are matched with $c_{j}$ and $c_{k}$. For each student $p_{k}$, her preference list contains every course in $C^{\prime}$ in increasing subscript order:
@@
The preference list of every $c_{j}\in C^{\prime}$ depends on the vertices adjacent to $u_{j}$. For every $w_{i}$ adjacent to $u_{j}$, $c_{j}$ includes in its preference list either $s_{i}^{1}$ or $s_{i}^{2}$, as well as $s_{i}^{3}$. Additionally, it includes every student in $P$. The order of all students is derived from the master list. If, for example, $u_{j}$ is adjacent to $w_{3}$, $w_{4}$ and $w_{5}$, the preference list of $c_{j}$ is given below. Here the superscript $l_{i}$ indicates a number that can be either $1$ or $2$. If $u_{j}$ is matched in $G$ then $c_{j}$ is matched with some agent from $S^{1}$ or $S^{2}$, otherwise it is matched with some $p_{k}$.
@@
The preference lists of every $d_{i}^{1}\in D^{1}$ and $d_{i}^{2}\in D^{2}$ are:
@@
Finally, the preference lists of $e_{1}$ and $e_{2}$ contain every student from $S^{3}$ and the preference list of $f$ contains every student from $S^{1}$ and $S^{2}$, in both cases in increasing subscript order. The students matched to $e_{1}$ and $e_{2}$ correspond to vertices in $G$ that are unmatched, while the students matched to $f$ correspond to vertices that are matched.
@@
Figure 7: Example of an Exact-MM instance, with $K=2$, and some of the preference lists in the transformed instance of ca-ml
@@
There are $n_{1}$ courses in $C^{\prime}$, with upper quota $1$ and number of credits $1$. There are also $n_{2}$ courses in each one of $D^{1}$ and in $D^{2}$, with upper quota $1$ and number of credits $2$. Finally, we have courses $e_{1}$ and $e_{2}$, with upper quota $n_{2}-K$ and number of credits $1$, and $f$, with upper quota $K$ and number of credits $1$. Hence, a course-complete matching has $n_{1}\cdot 1\cdot 1+2\cdot n_{2}\cdot 2\cdot 1+2\cdot(n_{2}-K)\cdot 1+1\cdot K\cdot 1=n_{1}+6n_{2}-K$ credits.
Likewise, there are $n_{2}$ students in each one of $S^{1}$, $S^{2}$ and $S^{3}$, with number of credits $2$, and $n_{1}-K$ students in $P$, with number of credits $1$. Thus, a student-complete matching has $3\cdot n_{2}\cdot 2+(n_{1}-K)\cdot 1=n_{1}+6n_{2}-K$ number of credits.
@@
We now show that a maximal matching $M$ of size $K$ in $G$ implies a student-complete (course-complete) pair-size-stable matching $M^{\prime}$ in $I$. For every pair $(u_{j},w_{i})\in M$, we add to $M^{\prime}$ either the pairs $(s_{i}^{1},c_{j})$ and $(s_{i}^{1},f)$, or the pairs $(s_{i}^{2},c_{j})$ and $(s_{i}^{2},f)$, depending on whether $s_{i}^{1}$ or $s_{i}^{2}$ has $c_{j}$ in her preference list. If $(s_{i}^{1},c_{j})\in M^{\prime}$, then we add $(s_{i}^{2},d_{i}^{1})$ to $M^{\prime}$, otherwise we add $(s_{i}^{1},d_{i}^{1})$. We also add $(s_{i}^{3},d_{i}^{2})$ to $M^{\prime}$. If $w_{i}$ has no partner in $M$, we add the following pairs to $M^{\prime}$: $(s_{i}^{1},d_{i}^{1})$, $(s_{i}^{2},d_{i}^{2})$, $(s_{i}^{3},e_{1})$, and $(s_{i}^{3},e_{2})$. Finally, let $j_{k}$ ($1\leq k\leq n_{1}-K$) be a sequence such that $w_{j_{k}}$ is unmatched in $M$. Then add $(p_{k},w_{j_{k}})$ to $M^{\prime}$. The matching $M^{\prime}$ is course-complete, as every course in $C^{\prime}$, $D^{1}$ and $D^{2}$ is matched with one student, $e_{1}$ and $e_{2}$ are matched with $n_{2}-K$ students, and $f$ is matched with $K$ students. It is also student-complete, because for every student $s\in S$, $T(s)=O(C_{M^{\prime}}(s))$. We show below that it is pair-size-stable.
@@
$M^{\prime}$ is a pair-size-stable matching.
@@
We prove the pair-size stability of $M^{\prime}$ by showing that no student can belong to a size-blocking pair. Every $s_{i}^{1}$ is matched with either $f$ and some $c_{j}$, which is her most preferred course, or with $d_{i}^{1}$. In either case, there is no course with which $s_{i}^{1}$ can form a size-blocking pair. Similarly, every $s_{i}^{2}$ is matched with either $f$ and some $c_{k}$, her most preferred course, with $d_{i}^{1}$, or with $d_{i}^{2}$. If $s_{i}^{2}$ is matched with $f$ and $c_{k}$, or with $d_{i}^{1}$, the same reasoning applies, and if she is matched with $d_{i}^{2}$, then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, so there is no size-blocking pair containing $s_{i}^{2}$.
We now consider $s_{i}^{3}$. This student is matched with either $d_{i}^{2}$ or with $e_{1}$ and $e_{2}$. In the first case, there is no course with an equal number of credits that $s_{i}^{3}$ prefers more, so there is no possible size-blocking pair. In the second case, $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$ and, because $M$ is maximal in $G$, $u_{j}$ and $u_{k}$ are matched with some $w_{l}$ and $w_{m}$. Thus, there exist $s_{l}^{1}$ (or $s_{l}^{2}$) and $s_{m}^{1}$ (or $s_{m}^{2}$) such that $c_{j}$ and $c_{k}$ are matched with them. Hence, $s_{i}^{3}$ cannot form a size-blocking pair with any other course. Finally, every $p_{k}\in P$ is matched with some $c_{j}\in C^{\prime}$. Let $c_{l}$ be a course that $p_{k}$ prefers over $c_{j}$. Then $c_{l}$ is matched with either some $s_{i}^{1}$ or $s_{i}^{2}$, or with a $p_{k}^{\prime}$ such that $p_{k}^{\prime}\succ_{MLS}p_{k}$. Thus, $p_{j}$ cannot form a size-blocking pair with $c_{l}$. ∎
Now, assume that there exists a course-complete (student-complete) pair-size-stable matching $M^{\prime}$ in $I$. We now show how to get a maximal matching $M$ of size $K$ in $G$. First, as $M^{\prime}$ is student-complete, every $p_{k}\in P$ is matched with some $c_{j}\in C^{\prime}$. As $M^{\prime}$ is course-complete, the $K$ remaining courses in $C^{\prime}$ are matched with students from $S^{1}$, $S^{2}$, or $S^{3}$. Similarly, as $M^{\prime}$ is course-complete, $|S_{M}(f)|=q^{+}(f)$, that is, course $f$ fills its capacity. Every student who has $f$ in her preference list belongs to either $S^{1}$ or $S^{2}$, and because $M^{\prime}$ is student-complete every student matched to $f$ must be matched to some $c_{j}$, so the $K$ courses from $C^{\prime}$ that are not matched with students from $P$ are matched with students from $S^{1}$ and $S^{2}$. We now show that for each subset $S_{i}\subseteq S$, either $s_{i}^{1}$ or $s_{i}^{2}$ is matched with some $c_{j}\in C^{\prime}$.
@@
If $s_{i}^{1}$ is matched with $c_{j}$, then $(s_{i}^{2},d_{i}^{1})$ must be in $M^{\prime}$, otherwise $d_{i}^{1}$ has no students and is not course-complete. Thus, $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$, or else $d_{i}^{2}$ has no students. If $s_{i}^{2}$ is matched with $c_{k}$, then, by the same reasoning, $(s_{i}^{2},d_{i}^{1})\in M^{\prime}$ and $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$.∎
Thus, we can build a matching $M$ in $G$ of size $K$ by taking every pair of the form $(s_{i}^{l},c_{j})\in M$, of which there are $K$, and adding the pair $(u_{j},w_{i})$ to $M$. For every $S_{i}$ such that no $s_{i}^{l}\in S_{i}$ is matched to some $c_{j}\in C^{\prime}$, we have that $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$, $(s_{i}^{3},e_{1})\in M^{\prime}$, and $(s_{i}^{3},e_{2})\in M^{\prime}$. We now show that $M$ is a maximal matching in $G$.
@@
Looking for a contradiction, assume that it is not, so that there exists unmatched $u_{j},w_{i}$ that are adjacent to each other. This implies that in $M^{\prime}$, $c_{j}$ is matched with some $p_{k}$ and $s_{i}^{3}$ is matched with $e_{1}$ and $e_{2}$. However, $s_{i}^{3}$ prefers $c_{j}$ over $e_{1}$ and $e_{2}$ and $c_{j}$ prefers $s_{i}^{3}$ over $p_{k}$. Thus, $(s_{i}^{3},c_{j})$ is a size-blocking pair, a contradiction to the stability of $M^{\prime}$. Therefore, $M$ is maximal. ∎
@@
We can modify this transformation to show that determining the existence of a course-complete (but not necessarily student-complete) pair-size-stable matching is NP-hard even if all conditions from the theorem statement hold and the length of the preference list of every student is bounded. We do so by removing every agent in $P$, and creating a new $p_{j}$ agent for every $c_{j}\in C^{\prime}$. The preference list of this $p_{j}$ contains $c_{j}$ only, and these $p_{j}$’s are at the end of the master list of students.

### v2

@@ 5.1 No lower quotas
For every $w_{i}\in W$, we create the following students: $s_{i}^{1}$, $s_{i}^{2}$, and $s_{i}^{3}$, with $T(s_{i}^{1})=T(s_{i}^{2})=T(s_{i}^{3})=2$. We also create courses $d_{i}^{1}$ and $d_{i}^{2}$, with $O(d_{i}^{1})=O(d_{i}^{2})=2$. Create the following subsets: $S_{i}=\{s_{i}^{1},s_{i}^{2},s_{i}^{3}\}$ ($1\leq i\leq n_{2}$), $D_{i}=\{d_{i}^{1},d_{i}^{2}\}$ ($1\leq i\leq n_{2}$), $S^{r}=\{s_{i}^{r}:1\leq i\leq n_{2}\}$ ($1\leq r\leq 3$), and $D^{r}=\{d_{i}^{r}:1\leq i\leq n_{2}\}$ ($1\leq r\leq 2$).
@@
For every $u_{j}\in U$, we create a course $c_{j}$, with $O(c_{j})=1$. Additionally, we create three courses $e_{1}$, $e_{2}$ and $f$, with $O(e_{1})=O(e_{2})=O(f)=1$, as well as $n_{1}-K$ students labelled $p_{1}$ to $p_{n_{1}-K}$, such that, for every $p_{k}$, $T(p_{k})=1$. The upper quota of every course is $1$, except for $e_{1}$ and $e_{2}$, which both have upper quota $n_{2}-K$, and $f$, which has upper quota $K$. Let $C^{\prime}=\{c_{j}:1\leq j\leq n_{1}\}$ and let $P=\{p_{k}:1\leq k\leq n_{1}-K\}$.
The master list of students ($\succ_{MLS})$ is
$s_{1}^{1}\succ_{MLS}s_{1}^{2}\succ_{MLS}s_{2}^{1}\succ_{MLS}s_{2}^{2}\succ_{MLS}\dots\succ_{MLS}s_{n_{2}}^{1}\succ_{MLS}s_{n_{2}}^{2}\succ_{MLS}\\ s_{1}^{3}\succ_{MLS}s_{2}^{3}\succ_{MLS}\dots\succ_{MLS}s_{n_{2}}^{3}\succ_{MLS}p_{1}\succ_{MLS}p_{2}\succ_{MLS}\dots\succ_{MLS}p_{n_{1}-K}$ (1)
@@
That is, the first students belong to subsets $S^{1}$ and $S^{2}$, such that every $s_{i}^{1}$ is followed by $s_{i}^{2}$. Then we place students from $S^{3}$, and finally students from $P$, in all cases in increasing subscript order. The master list of courses ($\succ_{MLC}$) is
$c_{1}\succ_{MLC}c_{2}\succ_{MLC}\dots\succ_{MLC}c_{n_{1}}\succ_{MLC}d_{1}^{1}\succ_{MLC}d_{1}^{2}\succ_{MLC}d_{2}^{1}\succ_{MLC}d_{2}^{2}\succ_{MLC}\dots\succ_{MLC}\\ d_{n_{2}}^{1}\succ_{MLC}d_{n^{2}}^{2}\succ_{MLC}e_{1}\succ_{MLC}e_{2}\succ_{MLC}f$ (2)
@@
If in the exact-mm instance $w_{i}$ is matched with $u_{j}$, then $s_{i}^{1}$ is to be matched with $c_{j}$ and $f$, $s_{i}^{2}$ with $d_{i}^{1}$ and $s_{i}^{3}$ with $d_{i}^{2}$. Similarly if $w_{i}$ is matched with $u_{k}$, then $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $c_{k}$ and $f$, and $s_{i}^{3}$ with $d_{i}^{2}$. If, however, $w_{i}$ is unmatched, $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{2}$, and $s_{i}^{3}$ with $e_{1}$ and $e_{2}$. Student $s_{i}^{3}$ contains courses $c_{j}$ and $c_{k}$ in her preference list to make sure that a course-complete first-coalition-stable matching in $I$ corresponds to a maximal matching in $G$, so that if $s_{i}^{3}$ is matched with $e_{1}$ and $e_{2}$, some students from $S^{1}$ or $S^{2}$ are matched with $c_{j}$ and $c_{k}$. For each student $p_{k}\in P$, her preference list contains every course in $C^{\prime}$ in increasing subscript order:
@@
The preference list of every $c_{j}\in C^{\prime}$ depends on the vertices adjacent to $u_{j}$. For every $w_{i}$ adjacent to $u_{j}$, $c_{j}$ includes in its preference list either $s_{i}^{1}$ or $s_{i}^{2}$, as well as $s_{i}^{3}$. Additionally, it includes every student in $P$. The order of all students is derived from $\succ_{MLS}$. If, for example, $u_{j}$ is adjacent to $w_{3}$, $w_{4}$ and $w_{5}$, the preference list of $c_{j}$ is given below. Here the superscript $l_{i}$ indicates a number that can be either $1$ or $2$. If $u_{j}$ is matched in $G$ then $c_{j}$ is matched with some student from $S^{1}$ or $S^{2}$, otherwise it is matched with some $p_{k}$.
@@
The preference lists of every course $d_{i}^{1}\in D^{1}$ and $d_{i}^{2}\in D^{2}$ are:
@@
Finally, the preference lists of courses $e_{1}$ and $e_{2}$ contain every student from $S^{3}$ and the preference list of course $f$ contains every student from $S^{1}$ and $S^{2}$, in all cases in increasing subscript order; in the case of $f$’s preference list, every $s_{i}^{1}$ is followed by $s_{i}^{2}$. The students matched to $e_{1}$ and $e_{2}$ correspond to vertices in $G$ that are unmatched, while the students matched to $f$ correspond to vertices that are matched.
@@
Figure 7: Example of an exact-$m$ instance, with $K=2$, and some of the preference lists in the transformed instance of ca-ml
@@
There are $n_{1}$ courses in $C^{\prime}$, each with upper quota $1$ and number of credits $1$. There are also $n_{2}$ courses in each of $D^{1}$ and in $D^{2}$, each with upper quota $1$ and number of credits $2$. Finally, we have courses $e_{1}$ and $e_{2}$, each with upper quota $n_{2}-K$ and number of credits $1$, and $f$, with upper quota $K$ and number of credits $1$. Hence, a course-complete matching has $n_{1}\cdot 1\cdot 1+2\cdot n_{2}\cdot 2\cdot 1+2\cdot(n_{2}-K)\cdot 1+1\cdot K\cdot 1=n_{1}+6n_{2}-K$ credits.
Likewise, there are $n_{2}$ students in each of $S^{1}$, $S^{2}$ and $S^{3}$, each with two available credits, and $n_{1}-K$ students in $P$, each with one available credit. Thus, a student-complete matching has $3\cdot n_{2}\cdot 2+(n_{1}-K)\cdot 1=n_{1}+6n_{2}-K$ number of credits.
@@
We now show that a maximal matching $M$ of size $K$ in $G$ implies a student-complete (course-complete by Lemma 7) pair-size-stable matching $M^{\prime}$ in $I$. For every pair $(u_{j},w_{i})\in M$, we add to $M^{\prime}$ either the pairs $(s_{i}^{1},c_{j})$ and $(s_{i}^{1},f)$, or the pairs $(s_{i}^{2},c_{j})$ and $(s_{i}^{2},f)$, depending on whether $s_{i}^{1}$ or $s_{i}^{2}$ has $c_{j}$ in her preference list. If $(s_{i}^{1},c_{j})\in M^{\prime}$, then we add $(s_{i}^{2},d_{i}^{1})$ to $M^{\prime}$, otherwise we add $(s_{i}^{1},d_{i}^{1})$ to $M^{\prime}$. We also add $(s_{i}^{3},d_{i}^{2})$ to $M^{\prime}$. If $w_{i}$ has no partner in $M$, we add the following pairs to $M^{\prime}$: $(s_{i}^{1},d_{i}^{1})$, $(s_{i}^{2},d_{i}^{2})$, $(s_{i}^{3},e_{1})$, and $(s_{i}^{3},e_{2})$. Finally, let $j_{k}$ ($1\leq k\leq n_{1}-K$) be a sequence such that $w_{j_{k}}$ is unmatched in $M$. Then add $(p_{k},c_{j_{k}})$ to $M^{\prime}$ ($1\leq k\leq n_{1}-K$). The matching $M^{\prime}$ is course-complete, as each course in $C^{\prime}$, $D^{1}$ and $D^{2}$ is matched with one student, each of $e_{1}$ and $e_{2}$ is matched with $n_{2}-K$ students, and $f$ is matched with $K$ students. It is also student-complete, because for every student $s\in S$, $T(s)=O(C_{M^{\prime}}(s))$. We show below that $M^{\prime}$ is pair-size-stable in $I$.
@@
$M^{\prime}$ is a pair-size-stable matching in $I$.
@@
We prove the pair-size stability of $M^{\prime}$ by showing that no student can belong to a size-blocking pair. Every $s_{i}^{1}$ is matched with either $f$ and some $c_{j}$, which is her most-preferred course, or with $d_{i}^{1}$. In either case, there is no course with which $s_{i}^{1}$ can form a size-blocking pair. Similarly, every $s_{i}^{2}$ is matched with either $f$ and some $c_{k}$, her most-preferred course, with $d_{i}^{1}$, or with $d_{i}^{2}$. If $s_{i}^{2}$ is matched with $f$ and $c_{k}$, or with $d_{i}^{1}$, the same reasoning applies as for $s_{i}^{1}$, and if she is matched with $d_{i}^{2}$, then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$. As $s_{i}^{1}\succ_{d_{i}^{1}}s_{i}^{2}$, there is no size-blocking pair containing $s_{i}^{2}$.
We now consider $s_{i}^{3}$. This student is matched with either $d_{i}^{2}$ or with $e_{1}$ and $e_{2}$. In the first case, there is no course with at least two credits that $s_{i}^{3}$ prefers to $d_{i}^{2}$, so there is no possible size-blocking pair. In the second case, $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$, $s_{i}^{2}\succ_{d_{i}^{2}}s_{i}^{3}$, and, because $M$ is maximal in $G$, $u_{j}$ and $u_{k}$ are matched in $M$ with some $w_{p}$ and $w_{q}$ respectively. Thus, there exist $s_{p}^{l_{p}},s_{q}^{l_{q}}\in S^{1}\cup S^{2}$ such that $(s_{p}^{l_{p}},c_{j})\in M^{\prime}$ and $(s_{q}^{l_{q}},c_{k})\in M^{\prime}$. Hence, $s_{i}^{3}$ cannot form a size-blocking pair with any other course. Finally, every $p_{k}\in P$ is matched with some $c_{j}\in C^{\prime}$. Let $c_{l}$ be a course that $p_{k}$ prefers over $c_{j}$. Then $c_{l}$ is matched with either some $s_{i}^{1}$ or $s_{i}^{2}$, or with a $p_{k}^{\prime}$ such that $p_{k}^{\prime}\succ_{MLS}p_{k}$. Thus, $p_{j}$ cannot form a size-blocking pair with $c_{l}$. ∎
Now, assume that there exists a course-complete (student-complete) pair-size-stable matching $M^{\prime}$ in $I$. We show how to construct a maximal matching $M$ of size $K$ in $G$. First, as $M^{\prime}$ is student-complete, every $p_{k}\in P$ is matched with some $c_{j}\in C^{\prime}$. As $M^{\prime}$ is course-complete, the $K$ remaining courses in $C^{\prime}$ are matched with students from $S^{1}$, $S^{2}$, or $S^{3}$. Similarly, as $M^{\prime}$ is course-complete, $|S_{M}(f)|=q^{+}(f)=K$, that is, course $f$ fills its capacity. Every student who has $f$ in her preference list belongs to either $S^{1}$ or $S^{2}$, and because $M^{\prime}$ is student-complete every student matched to $f$ must be matched to some $c_{j}$, so the $K$ courses from $C^{\prime}$ that are not matched with students from $P$ are matched with students from $S^{1}$ and $S^{2}$. We now show that for each subset $S_{i}$ ($1\leq i\leq n_{2}$), either $s_{i}^{1}$ or $s_{i}^{2}$ is matched with some $c_{j}\in C^{\prime}$, but not both.
@@
If $(s_{i}^{1},c_{j})\in M^{\prime}$, then $(s_{i}^{2},d_{i}^{1})$ must be in $M^{\prime}$, otherwise $d_{i}^{1}$ has no students and $M^{\prime}$ is not course-complete. Thus, $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$, or else $d_{i}^{2}$ has no students. If $(s_{i}^{2},c_{k})\in M^{\prime}$, then, by the same reasoning, $(s_{i}^{2},d_{i}^{1})\in M^{\prime}$ and $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$.∎
Thus, we can build a matching $M$ in $G$ of size $K$ by adding the pair $(u_{j},w_{i})$ to $M$ for every pair of the form $(s_{i}^{l_{i}},c_{j})\in M^{\prime}$, of which there are exactly $K$. For every $S_{i}$ such that no $s_{i}^{l_{i}}\in S_{i}$ is matched to some $c_{j}\in C^{\prime}$, we have that $\{(s_{i}^{1},d_{i}^{1}),(s_{i}^{2},d_{i}^{2}),(s_{i}^{3},e_{1}),(s_{i}^{3},e_{2})\}\subseteq M^{\prime}$. We now show that $M$ is a maximal matching in $G$.
@@
Looking for a contradiction, assume that $M$ is not maximal, so that there exists unmatched $u_{j},w_{i}$ that are adjacent to each other in $G$. This implies that in $M^{\prime}$, $c_{j}$ is matched with some $p_{k}$ and $s_{i}^{3}$ is matched with $e_{1}$ and $e_{2}$. However, $s_{i}^{3}$ prefers $c_{j}$ over each of $e_{1}$ and $e_{2}$, and $c_{j}$ prefers $s_{i}^{3}$ over $p_{k}$. Thus, $(s_{i}^{3},c_{j})$ is a size-blocking pair, a contradiction to the pair-size stability of $M^{\prime}$ in $I$. Therefore, $M$ is maximal in $G$. ∎
@@
We can modify this transformation to show that determining the existence of a course-complete (but not necessarily student-complete) pair-size-stable matching is NP-hard even if all conditions from the theorem statement hold and the length of the preference list of every student is bounded. We do so by removing every student in $P$, and creating a new $p_{j}$ student for every $c_{j}\in C^{\prime}$. The preference list of this $p_{j}$ contains $c_{j}$ only, and these $p_{j}$ students are at the end of the master list of students. We would then have that the maximum length of each student’s preference list is $5$.